A body starts from rest and reaches a velocity of 5m/s after travelling with uniform acceleration for 2s. calculate the accelaration


1
Expert's answer
2022-06-01T12:49:56-0400
"a=\frac{\Delta v}{\Delta t}=\frac{5-0}{2}=2.5\:\rm m/s^2"